Prove that limit x approach to 1 (28/(4x+3))=4
1
Expert's answer
2020-09-09T19:13:30-0400

We can prove it by epsilon and delta definition.

By definition, "\lim_{x\to a} f(x)=b" if and only if for every "\u03b5>0" there is a "\u03b4>0" such that: for all "x", if "0<|x\u2212a|<\u03b4", then "|f(x)\u2212b|<\u03b5".

As we want to show that "\lim_{x\to 1} \frac{28}{4x+3}=4", we have to make "|\frac{28}{4x+3}-4|<ε" and we control "|x-1|<\u03b4".

"|\frac{28}{4x+3}-4|=|\frac{28-4(4x+3)}{4}|=|\frac{28-16x-12}{4}|=|\frac{16-16x}{4}|=|\frac{16(1-x)}{4}|=|4(1-x)|".

In order to make "|4(x\u22121)|<\u03b5", it suffices to make "|x−1|<\frac{ε}{4}".

Proof. Let "\u03b5>0" be given, choose "δ=\frac{ε}{4}" and assume that "0 < | x \u2013 1| <\u03b4".

We have: if "0 < | x \u2013 1| <\u03b4" then

"|\frac{28}{4x+3}-4|=|4(1-x)|=4|1-x|=4|x-1|<4δ=4\frac{ε}{4}=ε".

Thus, we conclude that,  "|\frac{28}{4x+3}-4|<ε".

We have shown that for any positive ε, there is a positive δ such that for all "x", if "0<|x\u22121|<\u03b4", then "|\frac{28}{4x+3}-4|<ε".

So, by the definition of limit, we have "\lim_{x\to 1} \frac{28}{4x+3}=4".


The other way to prove it:

As the function "f(x)=\frac{28}{4x+3}" is continuous at point "x=1", then "\lim_{x\to 1} f(x)=f(1)=\frac{28}{4 \cdot 1+3}=\frac{28}{7}=4."